# Effective Quantization of Muon Optimizer States

### v1

@@ Abstract
The Muon optimizer, based on matrix orthogonalization, has recently shown faster convergence and up to 2× computational efficiency over AdamW in LLM pretraining. Like AdamW, Muon is stateful, requiring storage of both model weights and accumulated gradients. While 8-bit AdamW variants mitigate this overhead using blockwise quantization, they are typically stable only under dynamic quantization - which improves stability on linear quantization for extreme values. In this paper, we introduce the 8-bit Muon optimizer using blockwise quantization, supporting both linear and dynamic schemes. We demonstrate that 8-bit Muon maintains stability under both, while delivering $\sim$74% reduction in memory footprint compared to full-precision Muon. In extensive experiments, 8-bit Muon closely matches the performance of Muon while outperforming AdamW and 8-bit AdamW in pre-training a 1.6B model on 4B FineWeb tokens. It also shows competitive results when fine-tuning the Llama 3.2 3B model on post-training data. We also provide a theoretical perspective to help explain this robustness under quantization.
@@ 1 Introduction
Scaling laws for large language models (LLMs) (Kaplan et al., 2020; Hoffmann et al., 2022) indicate that larger models generally achieve better out-of-distribution performance across diverse tasks. Yet, GPU high-bandwidth memory (HBM) capacity has not kept pace with parameter counts. During training, memory is dominated by model parameters, gradients, optimizer states, and activations. Systems work has therefore focused on distributing these tensors across devices via distributed data parallel (DDP), Fully Sharded Data Parallel (FSDP) (Zhao et al., 2023), ZeRO stage-3 in DeepSpeed (Rajbhandari et al., 2020), and tensor/model parallelism (Shoeybi et al., 2019) in order to improve inference performance.
@@
Orthogonal to sharding is compressing the optimizer state. AdamW (Loshchilov & Hutter, 2017; Kingma, 2014), the de facto optimizer for LLMs, maintains two FP32 moment buffers (first and second moments) per parameter. For an 8B-parameter model (e.g., an 8B Llama-3 variant (Dubey et al., 2024)), this alone occupies $64$ GB ($\sim$80% of an NVIDIA H100’s 80 GB HBM), leaving little headroom for parameters, gradients, and activations. To mitigate this, Dettmers et al. (2021) quantize Adam’s optimizer states to 8 bits via block-wise dynamic (non-linear) quantization, preserving stability in the presence of extreme values while reducing optimizer memory by roughly $4\times$ — enabling performant training under tight memory budgets.
@@
Recently, there has been a surge of interest in moving beyond AdamW to improve training efficiency (Anil et al., 2020; Shazeer & Stern, 2018; Vyas et al., 2024). Among various advances, one particularly promising optimizer is Muon (Jordan et al., 2024) [^1] Muon is closely related to SGD with momentum, adding a per-layer matrix orthogonalization, which orthogonalizes the gradient momentum before updating the parameters. Equalizing the importance of all update directions results in improved stability and better convergence (Bernstein & Newhouse, 2024a; Bernstein, 2025). Several large-scale studies have confirmed Muon’s ability to achieve a 2x efficiency in a target validation loss compared to AdamW on a compute-optimal setup  (Liu et al., 2025; Shah et al., 2025). This includes extremely large models up to a trillion parameters, such as Kimi K2 (Team et al., 2025) and GLM4.5 (Zeng et al., 2025).
XL0551010151512.212.23.13.16.76.72.62.61.71.7Memory (GB)GPT-XLLlama-3.2-3B01010202024246.16.113.413.45.65.63.43.4Llama-3.2-3BAdamW-32AdamW-8DMuon-32Muon-8D/AdamW-32Muon-8D
Figure 1: Optimizer state memory (GB) for GPT-XL (1.6B) and Llama-3.2-3B. Muon-8D reduces optimizer state for the GPT-XL and Llama models by as much as 86% when compared to AdamW-32, and 74% when compared to Muon-32. Table 1 defines the different variants in the legend.
@@
In this paper, we introduce the 8-bit Muon optimizer with blockwise quantization. While 8-bit AdamW variants are usually stable only using dynamic quantization, we demonstrate that 8-bit Muon can handle both linear and dynamic quantization effectively. Our contributions can be summarized as:
  • We study 8-bit AdamW, and pinpoint the source of its instability with linear quantization. Surprisingly, we also discover that SGD with momentum [^2] Throughout, we use SGD to mean SGD with momentum, unless otherwise noted. and its closely related algorithm Muon are robust to linear quantization, achieving close to generalization parity when compared to their full-precision counterparts. We also provide a theoretical perspective to explain SGD’s and Muon’s robustness under quantization..
  • We propose and evaluate several 8-bit variants of the Muon optimizer, systematically applying both linear and dynamic quantization of Muon to hidden matrix-valued parameters and AdamW to the remaining parameters. Table 1 describes all the variants.
  • For pre-training models up to 1.6B parameters, our 8-bit Muon variants, regardless of the quantization scheme used for the Muon-associated parameters, are highly effective, achieving a validation loss within 1%-2% of the full-precision Muon optimizer for smaller models and matching its performance for the 1.6B model, while consistently outperforming all AdamW baselines.
  • For fine-tuning, we train the Llama 3.2 3B model (Dubey et al., 2024) on splits of the tulu-3-sft-mixture post-training dataset (Lambert et al., 2024). 8-bit Muon demonstrates competitive performance when compared to Muon, AdamW and 8-bit AdamW variants.
  • Our 8-bit Muon variants reduce the optimizer state memory substantially. For the 1.6B model, an 8-bit variant has a state that is smaller by up to $\sim$86% compared to AdamW (Muon-8D vs. AdamW-32), $\sim$74% compared to Muon (Muon-8D vs. Muon-32) and $\sim$44% compared to 8-bit AdamW (Muon-8D vs. AdamW-8D); enabling efficient training of large models. We achieve a similar reduction in optimizer state for SFTing the Llama 3.2 3B model, reducing the optimizer state footprint from $\sim$24 GB (AdamW-32) to $\sim$3.4 GB (Muon-8D) (Figure 1).
@@
8-bit Muon (dynamic), 8-bit AdamW 8b D 8b D Muon-8D
8-bit Muon (linear), 8-bit AdamW 8b L 8b L Muon-8L
@@
8-bit Muon (linear), 32-bit AdamW 8b L 32b Muon-8L/AdamW-32
Table 1: Optimizer variants considered for pre-training and SFT. D = dynamic quantization, L = linear quantization. 8-bit Muon has 4 variants.
@@ 2 Related Work
Efficient Optimizers There has been some prior work on alleviating the memory cost of training with optimizers like AdamW. Techniques like low-rank adaptation (LoRA) (Hu et al., 2022) allow a small subset of parameters to be fine-tuned for downstream tasks, but often fall behind in quality when compared to full parameter fine-tuning (Biderman et al., 2024). Adafactor (Shazeer & Stern, 2018) factorizes the second moment for matrices, reducing memory consumption compared to AdamW.  Dettmers et al. (2021) introduced the 8-bit Adam optimizer, but it only works when combined with careful blockwise + dynamic quantization. Adam Galore (Zhao et al., 2024) leverages the low-rank structure of the gradients to reduce state size and can be combined with state quantization. Our work is directly comparable to Dettmers et al. (2021)’s work, and can potentially be combined with low-rank updates.
Gradient Orthogonalization The Muon optimizer (Jordan et al., 2024) has sparked interest in algorithms that take advantage of the orthonormalization of the gradient of matrix-valued parameters. Muon makes each weight matrix update orthonormal through polar decomposition (Newton-Schulz), giving direction-only, spectrally controlled steps for hidden layers. To scale it to large LLMs, recent works add decoupled weight decay and careful per-parameter update scaling (Liu et al., 2025). Dion (Ahn et al., 2025) leverages orthonormalization but is built for distributed training: using low-rank orthonormalization with device-local momentum/error-feedback to avoid reconstruction or synchronization of full matrices.
@@ 3.1 Quantization for optimizers
Quantization is the process of reducing the precision of numerical representations by mapping a value expressed in a richer way into a simpler one. For example, representing a real number as a 32-bit floating-point value, or converting a 32-bit float into an 8-bit integer, are both forms of quantization. In deep learning, model parameters and optimizer states are typically stored as 32-bit floating-point numbers, making their conversion to lower-precision formats a primary goal of quantization.
For instance, linear quantization is a method (among many) to quantize the state tensor X of an optimizer from 32-bit floats to 8-bit integers. This involves (a) dividing all the floats in the tensor by the absolute max to get a normalization constant $S=\max(|\textbf{X}|)$ and (b) mapping the normalized value to the integer $i$ in the range 127-127 to $127$, using uniform spacing of the normalized values in [1,1][-1,1], written as $i=\text{round}\left(\frac{\textbf{X}}{S}\times 127\right)$. Dequantization then converts the codebook index back by multiplying with the previously stored normalization constant to recover $\widetilde{\mathbf{X}}=\frac{S}{127}\times i$.
@@
A crucial recipe for reducing the effect of outliers is blockwise quantization (Dettmers et al., 2021), where one can segment a tensor into one or more blocks, and then perform quantization (linear, dynamic, or other schemes) separately within each block. A nice side effect of blockwise quantization is that each block can be processed in parallel. For a more detailed treatment of blockwise and dynamic quantization, please refer to (Dettmers et al., 2021).
@@ 3.2 The Muon algorithm
$\mathbf{M}^{(t)}$ $=\beta\,\mathbf{M}^{(t-1)}+\nabla f_{t}\!\big(\mathbf{W}^{(t-1)}\big),$ (1)
$\mathbf{U}^{(t)}$ $=\mathrm{NS}\!\big(\mathbf{M}^{(t)}\big),$
$\mathbf{W}^{(t)}$ $=\mathbf{W}^{(t-1)}-\alpha\,\mathbf{U}^{(t)}.$
where $t$ is the iteration number and $\mathbf{M}$ is the momentum of the gradient. NS stands for the Newton-Schulz iteration process (Bernstein & Newhouse, 2024b; Higham, 2008), used to find an approximation for $\mathbf{U}\mathbf{V}^{T}$ where $\mathbf{U}\mathbf{\Sigma}\mathbf{V}^{T}$ is the singular value decomposition (SVD) of $\mathbf{M}$. Orthogonalization equalizes the importance of each update direction by collapsing all singular values to 1.
@@
$\mathbf{M}^{(t)}$ $=\beta\,\mathbf{M}^{(t-1)}+\nabla f_{t}\!\big(\mathbf{W}^{(t-1)}\big),$ (2)
$\mathbf{U}^{(t)}$ $=\mathrm{NS}\!\big(\mathbf{M}^{(t)}\big),$

### v3

@@ Abstract
The Muon optimizer, based on matrix orthogonalization, has recently shown faster convergence and better computational efficiency over AdamW in LLM pre-training. However, the memory overhead of maintaining high-precision optimizer states remains a challenge for large-scale deployment. In this paper, we introduce the 8-bit Muon optimizer using blockwise quantization.
In extensive Chinchilla-optimal experiments on pre-training models of up to 2.7B in size and fine-tuning them for instruction following, we demonstrate that 8-bit Muon achieves parity with Muon in terms of validation loss and downstream benchmarks, while achieving up to a 62% reduction in optimizer state footprint. Crucially, we show that Muon’s update mechanism is uniquely compatible with a simple linear quantization scheme, bypassing the complex dynamic scaling required for quantized AdamW. We supplement our empirical findings with a theoretical analysis of Muon’s robustness to quantization noise.
Machine Learning, Optimizer Quantization, Muon Optimizer
@@ 1 Introduction
Scaling laws for large language models (LLMs) (Kaplan et al., 2020; Hoffmann et al., 2022) indicate that larger models generally achieve better out-of-distribution performance across diverse tasks. Yet, GPU high-bandwidth memory (HBM) capacity has not kept pace with parameter counts. During training, memory is dominated by model parameters, gradients, optimizer states, and activations. Systems work has therefore focused on distributing these tensors across devices via distributed data parallel (DDP), Fully Sharded Data Parallel (FSDP) (Zhao et al., 2023), ZeRO stage-3 in DeepSpeed (Rajbhandari et al., 2020), and tensor/model parallelism (Shoeybi et al., 2019) in order to improve training performance.
@@
Orthogonal to sharding is compressing the optimizer state. AdamW (Loshchilov and Hutter, 2017; Kingma, 2014), the de facto optimizer for LLMs, maintains two FP32 moment buffers (first and second moments) per parameter. For an 8B-parameter model (e.g., an 8B Llama-3 variant (Dubey et al., 2024)), this alone occupies $64$ GB ($\sim$80% of an NVIDIA H100’s 80 GB HBM), leaving little headroom for parameters, gradients, and activations. To mitigate this, Dettmers et al. (2021) quantize Adam’s optimizer states to 8 bits via block-wise dynamic (non-linear) quantization, preserving stability in the presence of extreme values while reducing optimizer memory by roughly $4\times$, enabling performant training under tight memory budgets.
@@
Recently, there has been a surge of interest in moving beyond AdamW to improve training efficiency (Anil et al., 2020; Shazeer and Stern, 2018; Vyas et al., 2024). Among various advances, one particularly promising optimizer is Muon (Jordan and others, 2024) [^1] Muon is closely related to SGD with momentum, adding a per-layer matrix orthogonalization, which orthogonalizes the gradient momentum before updating the parameters. Equalizing the importance of all update directions results in improved stability and better convergence (Bernstein and Newhouse, 2024a; Bernstein, 2025). Several large-scale studies have confirmed Muon’s ability to achieve improved convergence rate to a target validation loss compared to AdamW on a compute-optimal setup  (Liu et al., 2025; Shah et al., 2025). Muon has also been used to train extremely large models up to a trillion parameters in size, like Kimi K2 (Team et al., 2025) and GLM4.5 (Zeng et al., 2025). While Muon offers improved convergence over AdamW, it still maintains FP32 optimizer state that competes for scarce HBM alongside parameters, gradients, and activations. As models continue to scale, this memory pressure remains a practical constraint even when the underlying optimizer improves.
1.1 Contributions
@@
In this paper, we introduce the 8-bit Muon optimizer with blockwise quantization. Whereas 8-bit AdamW variants are usually stable only with dynamic quantization, we demonstrate that 8-bit Muon can handle both linear and dynamic quantization effectively.
We systematically study approaches to quantizing the Muon optimizer by pretraining GPT models of up to 2.7B parameters with Chinchilla-optimal data ratios on the FineWeb-Edu set. Notably, we show that dynamic and linear quantization of Muon optimizer updates for pre-training workloads demonstrate parity in terms of quality - both for validation loss and downstream benchmarks - while achieving a 62% reduction in memory usage at the 2.7B model size. This memory reduction can be improved to 75% by additionally applying dynamic quantization to non-hidden-layer parameters updated by AdamW, at the cost of up to a 0.5% increase in validation loss. We observe similar results under Supervised Fine-Tuning (SFT), where memory efficiency is especially important in memory-constrained deployment settings.
A key empirical finding is that linear quantization of Muon optimizer state achieves parity with dynamic quantization and full-precision Muon in training progress. This contrasts with AdamW, where even sophisticated dynamic quantization does not fully retain performance, and linear quantization results in model divergence. This motivates our mechanistic analysis below.
Prior work by Dettmers et al. (2021) observed that naive 8-bit linear quantization can make AdamW unstable, and motivated block-wise dynamic quantization by highlighting sensitivity to quantization error in the optimizer states, especially the second-moment buffer. In Theorem 1, we formalize this mechanism by showing that AdamW can exhibit unbounded error under linear quantization, and we identify the second-moment accumulator in the denominator as the critical driver. This points to a natural control case: SGD with momentum, which removes the second-moment term. In Theorem 2, we show that linearly-quantized SGD+M admits bounded error at each optimizer step, and empirically verify that it remains effective in practice. Finally, in Theorem 3 we extend the same perspective to Muon, proving an analogous error bound with an additional dependence on the spectrum of the momentum matrix.
Key takeaway: Linear 8-bit quantization provides a simple and practical default to compress Muon’s optimizer state, achieving quality parity while substantially reducing optimizer state memory.
@@
8-bit Muon (linear), 8-bit AdamW (dynamic) 8b L 8b D Muon-8L/AdamW-8D
8-bit Muon (linear), 32-bit AdamW 8b L 32b Muon-8L/AdamW-32
8-bit Muon (dynamic), 8-bit AdamW (dynamic) 8b D 8b D Muon-8D/AdamW-8D
@@
Table 1: Optimizer variants considered for pre-training and SFT. D and L denote dynamic and linear quantization, respectively. The 8-bit Muon family comprises five studied variants. AdamW-8L related variants are unstable under Theorem 1, consistent with empirical observations, and hence omitted from the table.
@@ 2 Related Work
Efficient Optimizers There has been some prior work on alleviating the memory cost of training with optimizers like AdamW. Techniques like low-rank adaptation (LoRA) (Hu et al., 2022) allow a small subset of parameters to be fine-tuned for downstream tasks, but often fall behind in quality when compared to full parameter fine-tuning (Biderman et al., 2024). Adafactor (Shazeer and Stern, 2018) factorizes the second moment for matrices, reducing memory consumption compared to AdamW.  Dettmers et al. (2021) introduced the 8-bit Adam optimizer, but it only works when combined with careful blockwise + dynamic quantization. Adam GaLore (Zhao et al., 2024) leverages the low-rank structure of the gradients to reduce state size and can be combined with state quantization. Our work is directly comparable to Dettmers et al. (2021)’s work, and can potentially be combined with low-rank updates.
Gradient Orthogonalization The Muon optimizer (Jordan and others, 2024) has sparked interest in algorithms that take advantage of the orthonormalization of the gradient of matrix-valued parameters. Muon makes each weight matrix update orthonormal through polar decomposition (Newton-Schulz), giving direction-only, spectrally controlled steps for hidden layers. To scale it to large LLMs, recent works add decoupled weight decay and careful per-parameter update scaling (Liu et al., 2025). Dion (Ahn et al., 2025) leverages orthonormalization but is built for distributed training: using low-rank orthonormalization with device-local momentum/error-feedback to avoid reconstruction or synchronization of full matrices.
@@ 3.1 Quantization for Optimizers
Quantization is the process of reducing the precision of numerical representations by mapping a value in a larger set to one in a smaller discrete set. For example, representing a real number as a 32-bit floating-point value, or converting a 32-bit float into an 8-bit integer, are both forms of quantization. In deep learning, model parameters and optimizer states are typically stored as 32-bit floating-point numbers, making their quantization to lower-precision formats a primary goal for memory efficiency.
For instance, linear quantization is a method (among many) to quantize the state tensor $\mathbf{x}$ of an optimizer from 32-bit floats to 8-bit integers. Specifically, the $8$-bit integer linear quantization of $\mathbf{z}$ may be computed as:
$\mathbf{z}=\operatorname{round}\Big(\frac{127}{\max_{i}|\mathbf{x}_{i}|}\cdot\mathbf{x}\Big),$
where $\operatorname{round}$ denotes rounding to the nearest integer entry-wise. Note that all entries of $\mathbf{x}$ are integers in the range 127-127 to $127$ and hence may be exactly represented by an $8$-bit signed integer. This state $\mathbf{z}$ may then be de-quantized by $\tilde{\mathbf{x}}=(\max_{i}|\mathbf{x}_{i}|/127)\cdot\mathbf{z}$, and maintains the property that $|\tilde{\mathbf{x}}_{j}-\mathbf{x}_{j}|\leq\|\mathbf{x}\|_{\infty}/254$ for all $j$, guaranteeing bounded reconstruction error.
@@
A crucial recipe for reducing the effect of outliers is blockwise quantization (Dettmers et al., 2021), where one can segment the optimizer state into one or more blocks, and then perform quantization (linear, dynamic, or other schemes) separately within each block. A nice side effect of blockwise quantization is that each block can be processed in parallel. For a more detailed treatment of blockwise and dynamic quantization, please refer to (Dettmers et al., 2021).
@@ 3.2 The Muon Algorithm
$\mathbf{M}^{(t)}$ $=\beta\,\mathbf{M}^{(t-1)}+\nabla f\!\big(\mathbf{W}^{(t-1)}\big),$ (1)
$\mathbf{O}^{(t)}$ $=\mathrm{NS}\!\big(\mathbf{M}^{(t)}\big),$
$\mathbf{W}^{(t)}$ $=\mathbf{W}^{(t-1)}-\alpha\,\mathbf{O}^{(t)}.$
where $t$ is the iteration number and $\mathbf{M}$ is the momentum of the gradient. NS stands for the Newton-Schulz iteration process (Bernstein and Newhouse, 2024b; Higham, 2008), used to find an approximation for the polar factor $\mathbf{U}\mathbf{V}^{T}$ where $\mathbf{U}\mathbf{\Sigma}\mathbf{V}^{T}$ is the singular value decomposition (SVD) of $\mathbf{M}$. Orthogonalization equalizes the importance of each update direction by collapsing all singular values to 1.
@@
$\mathbf{M}^{(t)}$ $=\beta\,\mathbf{M}^{(t-1)}+\nabla f\!\big(\mathbf{W}^{(t-1)}\big),$ (2)
$\mathbf{O}^{(t)}$ $=\mathrm{NS}\!\big(\mathbf{M}^{(t)}\big),$